OpenStudy (anonymous):

How can you write the following expression in factorial notation? { ( n + 1 ) n ( n - 1 ) ( n - 2 )......( n - k + 1) } / ( n - 2 )

OpenStudy (rational):

try multiplying by \(\large \dfrac{(n-k)!}{(n-k)!}\)

OpenStudy (anonymous):

i try

OpenStudy (rational):

\[\large \dfrac{ ( n + 1 ) n ( n - 1 ) ( n - 2 )\cdots( n - k + 1)(n-k)! }{ (n - 2)(n-k)! }\]

OpenStudy (rational):

notice that top expression is exactly same as (n+1)! : \[\large \dfrac{ ( n + 1 ) ! }{ (n - 2)(n-k)! }\]

OpenStudy (rational):

you want to write it as a binomial coefficient is it ?

OpenStudy (anonymous):

thank you